find the equation of the line that passes through (-3.5) and perpendincular to a line 3x-6y-7 = 0 (p.s put this line in the form of y=mx+c)
Answer by stanbon(75887)

find the equation of the line that passes through (-3.5) and perpendincular to a line 3x-6y-7 = 0 (p.s put this line in the form of y=mx+c)
-----------------------------
If 3x-6y-7 = 0
6y = 3x-7
y = (1/2)x-(7/6)
Slope = 1/2
--------
Any line perpendicular to the given line must
have slope = -2
-----------------------
Find equation of line thru (-3,5) with slope = -2.
Form: y = mx+b
5 = -2*-3+b
b = -1
----
Equation:
y = (1/2)x-1
==================
Cheers,
Stan H.
================